OpenStudy (anonymous):

using the following equation, find the center and radius of the circle. x^2+2x+y^2+4y=20 medal and fan will be given

OpenStudy (anonymous):

@IMStuck hi can you help me please

OpenStudy (anonymous):

@ganeshie8 hi can you help me please

OpenStudy (anonymous):

@phi can you help me please

OpenStudy (anonymous):

@mathmale hey can you help me with this please

OpenStudy (anonymous):

@mathstudent55 hi can you help me please

OpenStudy (mathmale):

Given the equation x^2+2x+y^2+4y=20, we must put it into the form of the standard equation of a circle with center (h,k) and radius r:\[(x-h)^2+(y-k)^2=r^2\] and so ...

OpenStudy (anonymous):

yess

OpenStudy (mathmale):

Starting with the given x^2+2x+y^2+4y=20, complete the square twice, first that of x^2+2x and next of y^2+4y. Have you done "completing the square" before?

OpenStudy (anonymous):

no ihavent

OpenStudy (mathmale):

I'll guide you thru completing the square for x^2+2x: The coefficient of 'x" is '2', right? Take half of this '2' Square the result. Do this now, please.

OpenStudy (anonymous):

would it be one?

OpenStudy (anonymous):

and isn't there a simpler way of doing this

OpenStudy (anonymous):

because i did something similar to this in trig

OpenStudy (mathmale):

Yes. Half of 2 is 1; the square of 1 is 1. Yes. Add this to x^2 + 2x, and then subtract the same quantity. Do this now, please. No...there is no easier way. ;)

OpenStudy (anonymous):

lol ok... and 1+2 =3

OpenStudy (anonymous):

i was supposed to substitute x for 1 right?

OpenStudy (mathmale):

You have x^2 + 2x. You identify the coefficient of x (which happens to be 2); you halve that, and square the result (obtaining 1), and then you add and immediately subtract that 1 from x^2 + 2x: x^2 + 2x + 1 - 1. Where did your '2' come from?

OpenStudy (mathmale):

gabs?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

I'm trying to read it again so I can understand

OpenStudy (anonymous):

wait one second I'm getting my notes

OpenStudy (mathmale):

Need quicker responses from you, please. Re write x^2 + 2x as x^2 + 2x + 1 -1 (which is exactly equivalent). then your problem becomes x^2 + 2x + 1 - 1 + y^2 +4y = 20

OpenStudy (anonymous):

i can't find them

OpenStudy (mathmale):

Please complete the square of y^2 + 4y in exactly the same fashion. Your expression? \[x^2+2x+1-1+y^2+4y +?-?=20\]

OpenStudy (mathmale):

Please finish this.

OpenStudy (mathmale):

Hint: what is the coefficient of y?

OpenStudy (mathmale):

Gabs? I need to know that you're there and are about to type something. If this is a bad time for you, perhaps we could finish later.

OpenStudy (anonymous):

I'm reading sorry

OpenStudy (anonymous):

coefficient of y is 4

OpenStudy (mathmale):

and half of that is what?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so you subtract 2 from the equation right

OpenStudy (mathmale):

And the square of 2 is what?

OpenStudy (mathmale):

No, you don't subtract 2 from the equation.

OpenStudy (mathmale):

the square of 2 is ... ?

OpenStudy (anonymous):

4

OpenStudy (mathmale):

So, you have y^2 + 4y and must now add 4 to that and then immediately subtract 4. What do you ob tain from following these steps?

OpenStudy (anonymous):

x and y

OpenStudy (mathmale):

Actually, your y^2 + 4y becomes y^2 + 4y + 4 - 4, doesn't it? We did the same thing with x^2 + 2x; we got x^2 +2x + 1 - 1, didn't we?

OpenStudy (anonymous):

yes and then what do we do?

OpenStudy (mathmale):

Rewrite the whole equation: x^2 + 2x + 1 - 1 + y^2 + (you finish it, please)

OpenStudy (anonymous):

4x-4

OpenStudy (mathmale):

Don't you mean 4y + 4 - 4? leading to x^2 + 2x + 1 - 1 + y^2 + +4y + 4 - 4 = 20 ?

OpenStudy (mathmale):

Lot of detail here, but it's essential that you undrstand these procedures.

OpenStudy (mathmale):

Note, please, that x^2 + 2x + 1 is the square of (x+1)^2, so we can re-write x^2 + 2x + 1 - 1 + y^2 + +4y + 4 - 4 = 20 as (x+2)^2 -1 + y^2 + +4y + 4 - 4 = 20

OpenStudy (mathmale):

Can you finish this? Rewrite y^2 + 4y + 4 as what perfect square?

OpenStudy (anonymous):

ok .... how is is this going to find the radius

OpenStudy (anonymous):

ill finish this thanks

OpenStudy (anonymous):

for your patience

OpenStudy (mathmale):

You should end up with (x+1)^2 + (y+2)^2 -1 -2 = 20 -1 -2 = 17. Compare this to \[(x-h)^2+(y-k)^2=r^2\] Thru this comparison, can you identify the value of 'h'?

OpenStudy (anonymous):

1 ?

OpenStudy (mathmale):

Actually, it'd be -1. Here's why: (x-h)^2 = (x+1)^2 So -h=+1, or +h=-1. OK?

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

:)

OpenStudy (mathmale):

You're welcome, gabs! Good luck!

OpenStudy (anonymous):

@phi can you help me please i didn't understand with the other person

OpenStudy (anonymous):

@thomaster hey can you help me

OpenStudy (phi):

how far did you get ?

OpenStudy (anonymous):

i honestly didn't understand anything he said

OpenStudy (anonymous):

i just need to find the radius and center... and did he get it right so far ? @phi

OpenStudy (phi):

I am sure he did. To do this, you use something called "completing the square" ring any bells.?

OpenStudy (anonymous):

no i was never taught this

OpenStudy (anonymous):

actually i understand so far but what about the rest? @phi

OpenStudy (phi):

start with x^2+2x+y^2+4y=20 add 1 to both sides x^2 + 2x + 1 + y^2 +4y = 20+1 add 4 to both sides (x^2 + 2x + 1) + (y^2 +4y +4) = 20+1+4 by algebraic magic we have (x^2 + 2x + 1) = (x+1)^2 and (y^2 +4y +4) = (y+2)^2 and on the right side 20+1+4= 25 we get (x+1)^2 + (y+2)^2 = 5^2

OpenStudy (anonymous):

and the radius would be 5^2

OpenStudy (anonymous):

and the center 1,2 right

OpenStudy (phi):

we want to match that up with \[ (x-h)^2 + (y-h)^2 = r^2\] we re-write x+1 as x- (-1) and y+2 as y - (-2) \[ (x - (-1))^2 + (y - (-2))^2 = 5^2 \]

OpenStudy (phi):

you have the right idea, but the details matter. 5^2 "matches up" with r^2 so r is not 5^2. r is 5

OpenStudy (phi):

and the center is -1,-2

OpenStudy (anonymous):

thank you soooo much you explained it much better :)